OpenStudy (anonymous):

361/9 a^2 = a^2 + 16 Show how you get 3 square root of 22/22 by solving for a

OpenStudy (cwrw238):

cross multiplying: 9a^4 + 144a^2 - 361 = 0 ley a^2 = A 9A^2 + 144A - 361 = 0 solve using quadratic formula them a^2 = roots of this equatiuon

OpenStudy (anonymous):

i don't understand what you mean by cross mulitiplying... and how did you get 144?

OpenStudy (cwrw238):

rewriting: 361 a^2 + 16 ---- = --------- 9 a^2 1 multiply top by bottom starting from 361 and bottom to top by 9 a^2 ( like a cross) so 361 * 1 = 9a^2 * (a^2 + 16)

OpenStudy (cwrw238):

so 361 = 9a^2 * a ^2 + 9^a2 * 16 361 = 9 a^4 + 9*16 a^2 361 = 9a^4 + 144 a^2

OpenStudy (cwrw238):

and writing in standard form 9a^4 + 144 a^2 - 361 = 0

OpenStudy (anonymous):

it's either 3square root of 22/22 or 12/square root of 352

OpenStudy (anonymous):

is it this? 

OpenStudy (anonymous):

yes, but I want to see the work.

OpenStudy (cwrw238):

in any case a^2 = - 144 +- sqrt((144^2 - 4*9*-361) ------------------------------ 18 = -8 +- sqrt(20736 + 12996) ------------------ 18 = -8 + 183.66 or -8 - 183.66 ------ ------ - this is discounted as a^2 cannot be negative 18 18 a^2 = -8 + 10.2 = 2.2 a = + sqrt 2.2 or - sqrt 2.2 = 1.48 or -1.48 which is different from your answers - of course i might be wrong!!

OpenStudy (cwrw238):

please clarify the left hand side of the equation is it (361/9) time s a^2 OR 361 divided by 9a^2 ?

OpenStudy (anonymous):

(361/9) times a^2

OpenStudy (cwrw238):

ah - thats different - it looked like my second option

OpenStudy (cwrw238):

u should have written it as ( 361 a^2) / 9 to avoid confusion so now its a different story..

OpenStudy (cwrw238):

( 361 a^2) / 9 = a^2 + 16 ( 361 a^2) / 9 - a^2 = 16 352 a^2 / 9 = 16 352 a^2 = 144 a = sqrt (144/352) = 12 / sqrt352 or -12 / sqrt352

OpenStudy (cwrw238):

at last!!

OpenStudy (anonymous):

THANKS SO MUCH !!

OpenStudy (cwrw238):

lol - yw

OpenStudy (anonymous):

sorry for wasting your precious time to help me. you're the best

OpenStudy (cwrw238):

thats ok